# Cosmological Simulations of Stellar Halos with Gaia Sausage–Enceladus Analogues: Two Sausages, One Bun?

### v1

@@ 3.3 Progenitors of GSE-like Debris
Generally, RA mergers are more recent ($t_{\mathrm{infall}}=5.9_{-2.0}^{+3.3}$ Gyr ago) and larger ($\log M_{\mathrm{dyn}}/\mathrm{M}_{\odot}=11.0^{+0.3}_{-0.1}$) than mergers contributing to RA pairs. Of those that contribute to RA pairs, the merger with the larger $f$ is typically more recent ($t_{\mathrm{infall}}=9.4^{+2.4}_{-2.7}$ versus $11.1^{+1.1}_{-2.4}$ Gyr ago) and has a slightly larger mass ($\log M_{\mathrm{dyn}}/\mathrm{M}_{\odot}=10.7^{+0.2}_{-0.4}$ versus $10.6^{+0.2}_{-0.7}$). This systematic shift in merger populations is particularly manifest in their star formation histories, explored in §5, but we first characterize the kinematic properties of these GSE analogues.
@@
*4
4 Kinematic Features
As summarized in Table 2, a number of GSE-like events are identified in TNG’s MW analogues. In general, RA debris is created by a single merger or—in a non-trivial fraction of cases—by pairs of mergers. A natural question that follows is whether or not these two scenarios can be distinguished using observable properties of the stellar halos. To this end, this section explores the kinematic properties of the stars in the RA mergers and pairs, such as their velocities, orbital energies and angular momenta. §5 will focus instead on stellar ages and abundances.
Figure 4 shows the distribution of spherical galactocentric velocities $v_{r}$ versus $v_{\phi}$ for the RA debris of four MW analogues. The star counts are indicated by the grayscale histogram, while the contours enclose 68% of the stars from each merger: the single RA merger in red and the large-$f$ (small-$f$) merger of the RA pair in purple (pink). The four panels illustrate the kinematic diversity across different MWs in the TNG set. The top row of Figure 4 provides two examples of single RA mergers. In the left panel, the debris has a clear “sausage”-like velocity distribution, with a distinctive extension along $v_{r}$. In the right panel, the RA merger shows a greater population of stars on less eccentric orbits, with $v_{r}\sim 0$ and nonzero $v_{\phi}$, making the full extent of the debris potentially more difficult to distinguish from older contributions to the ex situ stellar halo that are expected to be nearly isotropic (Necib et al., 2019b).
Typically, single RA mergers have lower $v_{\phi}$ dispersion than RA pairs: the former have $\sigma_{\phi}=86^{+13}_{-18}~{}\mathrm{km\ s}^{-1}$ while the latter have $\sigma_{\phi}=95^{+14}_{-11}~{}\mathrm{km\ s}^{-1}$. However, because there is significant scatter in the azimuthal dispersion from halo to halo, the precise value of $\sigma_{\phi}$ does little to distinguish an individual RA merger from a pair. The other components of velocity are more comparable within their scatter. In the radial direction, single RA mergers exhibit a dispersion of $\sigma_{r}=159_{-27}^{+18}~{}\mathrm{km\ s}^{-1}$ versus the pairs’ $164_{-21}^{+29}~{}\mathrm{km\ s}^{-1}$. In the polar direction, RA mergers have $\sigma_{\theta}=90_{-20}^{+12}~{}\mathrm{km\ s}^{-1}$ versus the pairs’ $89_{-5}^{+12}~{}\mathrm{km\ s}^{-1}$. These trends (in addition to the trends in $\beta$ shown in the right panel of Figure 2) do indicate that debris from the RA pairs is slightly more isotropized, befitting of their earlier infall times (cf. Figure 3), though this is only evident in their $v_{\phi}$ values.
The debris from the same MWs are shown in Figure 5, where the distribution of orbital energy $E$ is plotted against $L_{z}$, the $z$-component of angular momentum. [^10] Because the shape of the distribution is more important than the numerical values of $E$ and $L_{z}$, the axes are rescaled. In particular, the energy is centered at $\bar{E}$, the median energy of the RA debris, and normalized by $\Delta E$, the difference between the 2nd and 98th percentile. Similarly, though $L_{z}$ is still centered at zero, it is normalized by $\Delta L_{z}$, the difference between the 2nd and 98th percentile values of RA debris’ $L_{z}$. As in Figure 4, the top row contrasts two examples of single RA mergers. The example on the right has more complex substructure, with debris deposited at distinct energies. On their own, distinct clumps in energy and angular momentum could potentially be misinterpreted as originating from multiple mergers—indeed, due to the inherent difficulties in effectively clustering stars in observational parameters, a standard approach is to accept this fact and intentionally over-fragment with clustering algorithms, so that multiple clusters of stars can later be associated to the same object (e.g. Koppelman et al., 2019; Yuan et al., 2020; Lövdal et al., 2022; Ou et al., 2023). These analyses must therefore be performed with great care, keeping in mind the potential complexities of stellar debris in these kinematic spaces.
Figure 6: Star formation histories for the ex situ stars of four MW analogues with RA debris. In each panel, the black histogram shows the total ex situ star formation, while the colored lines show the contribution from individual mergers, with single RA mergers in red, large-$f$ components of RA pairs in purple, and small-$f$ components of RA pairs in pink. The vertical dashed line shows the infall time, $t_{\mathrm{infall}}$, of each merger, and the dotted line shows the quenching time, $t_{90}$, by which 90% of stars have formed. The left column highlights cases in which star formation ceases shortly after the galaxies enter the virial radius of their host. In the case of the RA pair (which is given a marker consistent with previous figures), the periods of star formation of the two galaxies comprising the pair are well-separated in time, making them easier to distinguish from one another. The right column highlights examples where star formation continues long after the galaxies are accreted onto their hosts and where the RA pair does not show the separation seen in the left panel. Versions of this figure for each choice of RA debris are included in Appendix C.
Figure 7: The quenching time $t_{90}$ relative to the infall time $t_{\mathrm{infall}}$, shown against the distance of first pericenter for each galaxy contributing to RA debris. All markers are consistent with Figure 3. Galaxies which continue forming stars after crossing the virial radius are below the gray $t_{90}=t_{\mathrm{infall}}{}$ line. While the quenching time is typically within a gigayear of the infall time, it is possible for star formation to continue well after the infalling galaxy crosses the virial radius. This is especially true of galaxies on orbits with large pericentric distances: in these cases, star formation can continue for many gigayears after infall.
Figure 8: Tinsley–Wallerstein diagram showing [Mg/Fe] abundance relative to [Fe/H] abundance for the same RA debris shown in Figure 6. Formatting is consistent with Figure 4 and Figure 5. A gold contour is also included for the in situ disk in each MW analogue to guide the eye. RA debris is generally metal-poor and $\alpha$-enhanced relative to the disk. This chemistry can in some cases distinguish the two populations of stars comprising an RA pair, e.g. in the bottom-left panel, which displays a clear $\alpha$-enhancement in the high-$f$ merger relative to the low-$f$ merger. However, this is not always the case; in the right column, the galaxies comprising the pair formed their stars at the same time, and the chemical abundances prove ineffective at separating the two. Versions of this figure for each choice of RA debris are shown in Appendix C.
Figure 9: Comparison of quantities relevant to star formation in each of the RA debris mergers. In each panel, the highlighted region runs from the 16th to the 84th percentile for the relevant quantity, with the scatter point at the 50th percentile (with markers consistent with previous figures). The Subfind ID of each MW analogue is also provided. (Left:) Formation time, $t_{\mathrm{form}}$, of stars from each merger relative to the merger’s infall time, including a colored vertical bar indicating $t_{90}$, the time at which 90% of the galaxy’s stars had formed. Generally, star formation continues for some time prior to infall, at which point it ceases, and $t_{90}\sim t_{\mathrm{infall}}$. As RA pairs tend to have earlier infall times and smaller masses (both in $M_{\mathrm{dyn}}$ and $M_{\star}$), their star formation is brief and does not extend as far prior to $t_{\mathrm{infall}}$ as it does for RA mergers. Furthermore, RA mergers may hold on to their gas more effectively and continue forming stars many gigayears after $t_{\mathrm{infall}}$. (Center:) Abundance of iron relative to hydrogen. (Right:) Abundance of magnesium relative to iron. These right two panels show some systematic trends among the RA debris, with RA pairs more metal poor and $\alpha$-enhanced than the single RA mergers, though these differences are less pronounced relative to the overall scatter in chemical abundances.
The bottom row of Figure 5 shows the RA pair examples in orbital energy and angular momentum space. In the left panel, the contributions from each merger are more easily distinguished from each other: only 12% of the high-$f$ merger stars fall within the pink contour and 19% of the low-$f$ stars fall within the purple contour. Conversely, the right panel’s RA pair is not easily distinguished as two independent mergers, as 36% of the high-$f$ merger stars fall within the pink contour and 50% of the low-$f$ fall within the purple contour. This overlap occurs even though the two mergers comprising the RA pair have infall times separated by more than a gigayear (Table 2) and have totally independent orbital initial conditions; their debris still settles into the same region of this kinematic space. There are hints that such overlap occurs in observational data, making it difficult to determine whether, e.g., structures like Sequoia, Arjuna, LMS-1, and I’itoi are truly independent structures or whether they should be considered part of the GSE debris (Massari et al., 2019; Koppelman et al., 2019; Horta et al., 2023).
[…]
*4
5 Star Formation Histories
Chemodynamical space is particularly effective at separating various populations of ex situ material, as stars from different infalling dwarf galaxies will generically have different formation environments and therefore different chemical abundances. Chemical information can also be used to infer the ages of ex situ stars, which gives insight into the quenching time and overall star formation history (SFH) of the dwarf in which they originated. Because there are systematic shifts in the infall times of single mergers and pairs of mergers which contribute RA debris (cf. §3.3), studying the SFH—and thence inferring the infall time of the dwarf galaxy progenitor(s)—is a promising route to distinguishing the two scenarios.
Infalling dwarfs are thought to stop forming stars quickly, as the MW’s circumgalactic medium strips them of their gas (Simpson et al., 2018; Akins et al., 2021; Engler et al., 2023). A close look at the satellites of MW-like hosts in TNG highlights the diversity of SFHs that can result from this process. In some cases—especially for the smallest, earliest-infalling satellites—star formation ceases soon after the satellite crosses the host’s virial radius. In other cases, the gas is not immediately stripped, and star formation continues for some time: for example, Engler et al. (2023) have shown that satellite quenching around MW analogues occurs 2.5 Gyr after infall for the median satellite in the TNG model and 7.5 Gyr after infall for the 84th-percentile satellite. Figure 6 illustrates the diversity across four example MWs in our sample. The horizontal axis corresponds to the formation time of ex situ stars, $t_{\rm form}$. The black histogram is the total SFH, while the colored histograms are the contributions from the RA debris. The infall time of each merger, $t_{\mathrm{infall}}$, is indicated by a dashed vertical line of the same color. Each dotted vertical line corresponds to the quenching time, $t_{90}$, by which 90% of a dwarf’s stars have formed.
The left column of Figure 6 showcases galaxies in which star formation ceases quickly after infall, with $t_{\mathrm{infall}}{}\sim t_{90}$. Furthermore, the galaxies comprising the RA pair (bottom-left panel) are well-separated in terms of the formation times of their stars, meaning the SFHs of these two galaxies are sufficient to distinguish them from each other. The right column, however, highlights cases in which this is not the case. For the single merger, in the top-right panel, star formation continues for some time following the dwarf’s infall, making it difficult to infer the $t_{\mathrm{infall}}$ from the SFH. For the RA pair, in the lower-right panel, the two galaxies form stars at very similar times, with overlapping SFHs which would make distinguishing the two stellar populations difficult on the basis of their age.
Whether or not star formation continues after $t_{\mathrm{infall}}$ is influenced by the galaxy’s orbit as it is accreted onto the MW analogue. In particular, Figure 7 shows that continued star formation is correlated with large pericentric distances: the horizontal axis shows the distance of the infalling galaxy’s first pericenter, and the vertical axis shows the quenching time $t_{90}$ relative to the infall time $t_{\mathrm{infall}}$. Most galaxies—including those shown in Figure 6, save the upper-right panel—cease forming stars within a gigayear of their infall, but this is especially true for galaxies on orbits with small pericenters. In these cases, the galaxy falls directly toward the MW analogue’s dense inner region and star formation is prevented due to gas loss and tidal stripping. Galaxies with larger pericentric distances—such as ID 520885, shown in the upper-right panel of Figure 6, which has a pericentric distance of 77 kpc—are less affected by stripping and are able to continue forming stars for many gigayears after their infall (cf. Di Cintio et al., 2021).
Including chemical abundances, especially for $\alpha$-process elements, reveals more detailed information about the environment in which the stars formed. Figure 8 shows a Tinsley–Wallerstein diagram for the same halos as Figure 6, with formatting similar to previous figures: the grayscale histograms show the RA debris, with contours enclosing 68% of the stars from each merger. For reference, a gold contour encloses in situ disk stars, i.e. stars bound to the MW analogue at the time of their formation which have orbital circularity $j_{z}/j_{\mathrm{circ}}>0.7$ (Scannapieco et al., 2009; Sotillo-Ramos et al., 2022). [^11] Here, $j_{z}$ is the $z$-component of the star’s specific angular momentum about the galactic center, and $j_{\mathrm{circ}}$ is the specific angular momentum of a circular orbit at the star’s current galactocentric distance.
While specific values of chemical abundances depend on the stellar physics model of TNG, [^12] The chemical abundances are in keeping with the current state-of-the-art, but they are based on metal yields which are rather uncertain (Pillepich et al., 2018). For example, it is known that TNG MW analogues are fairly $\alpha$-enriched relative to observations, cf. Naiman et al. (2018). Semenov et al. (2024, §2.3) provide approaches to correcting the metallicity distributions of TNG MWs, but in this work we instead focus on systematic shifts. relative shifts from one merger to another (and of mergers relative to the disk) still serve as an interesting comparison point. In the bottom row of Figure 8, the left panel shows the RA pair comprised of mergers with more distinct SFHs. Here, the two components have a clear separation in [Mg/Fe] relative to each other, with two sequences present in the histogram. In this pair, 10% of the high-$f$ merger stars fall within the pink contour and 2% of the low-$f$ merger stars fall within the purple contour. The lower-right panel shows the example RA pair with concurrent SFHs, and here the $\alpha$ abundance is not enough to clearly distinguish the two mergers from each other. In this pair, 54% of the high-$f$ merger stars fall within the pink contour and 39% of the low-$f$ merger stars fall within the purple contour.
Figure 9 summarizes the SFHs and chemical abundances across the sample: for each galaxy contributing to RA debris, a marker and horizontal line show the median and spread of the relevant quantity. The left panel shows the duration of star formation with respect to the infall time of the mergers, with an additional marker for the quenching time, $t_{90}$, at which 90% of the merger’s stars are formed. The duration of star formation is one suggestive distinction between RA mergers (in red) and RA pairs (in purple and pink), as the former have SFHs which extend well before their $t_{\mathrm{infall}}$, while the latter form their stars quickly, as they are generally accreted earlier in cosmic history (cf. Figure 3). The panel also reveals that $t_{90}$ generally serves as a good proxy for $t_{\mathrm{infall}}$; for most galaxies in this sample, the $t_{90}$ marker is close to the line through the origin which denotes the infall time. Notably, the exceptions to this trend are often single RA mergers, rather than the mergers of a RA pair; in terms of inferring $t_{\mathrm{infall}}$ from $t_{90}$, this further biases RA mergers toward later-inferred infall times.
The clear differences in SFH are less concretely reflected in chemical abundances in the TNG model, though two systematic shifts are still present. First, in [Fe/H], RA pairs are biased in the metal-poor direction relative to single RA mergers, indicative of the early epochs in which their stars formed. Second, in [Mg/Fe], the RA pairs are biased toward $\alpha$-enhancement, especially for the smaller-$f$ component of the pair (shown in pink). While these trends may not be sufficient to distinguish individual mergers from pairs (or galaxies comprising pairs from each other), the existence of the systematic biases between populations serves as a potential indicator for the merger histories of MW analogues. This is especially true if stellar ages can be robustly determined, as these trends are more apparent in SFH than in stellar chemistry.
[…]
*4
6 Conclusion
This work presents a sample of 98 MW analogues in the TNG50 simulation of the IllustrisTNG suite. Each analogue has its full accretion history identified with each ex situ star assigned to the merger that contributed it (§2.3). This is the largest sample of MW analogues for which this decomposition has been performed, and it allows for unprecedented statistics in studying the assembly of the MW and the buildup of its stellar halo (§3.1).
The MW analogues are chosen to match observations of the Galaxy’s stellar mass and local environment (§2.2), though certain aspects of the MW itself are rare across the sample. In particular, no analogues have an LMC-like satellite with mass $M_{\mathrm{dyn}}>10^{11}~{}\mathrm{M}_{\odot}$, indicating that the presence of the LMC distinguishes our Galaxy as an unlikely system. Further, while 33% of MW analogues have stars with properties similar to the GSE debris seen in our Galaxy—namely, highly radially-biased velocities comprising a majority of the inner ex situ stars (§3.2)—only 4% of MWs have debris as anisotropic as the GSE debris. This is in line with other studies which find the MW to be rare: see, e.g., Evans et al. (2020); Buch et al. (2024).
With the large sample of MW analogues, we can uncover the origins of GSE-like debris in a $\Lambda$CDM context. In particular, we consider scenarios in which the GSE debris is comprised of multiple independent mergers, rather than being accreted all at once. Approximately a third of the GSE-like debris seen in TNG is built up from two mergers. Statistics alone, therefore, do not preclude the possibility that the GSE debris in our own Galaxy is comprised of such a pair of mergers.
@@ 6 Conclusion
Given the possibility of a two-merger GSE, the primary concern is in distinguishing it from a single-merger origin; however, it proves difficult for observable parameters to differentiate the two scenarios. Kinematics alone provide little conclusive evidence, as shown in §4. Single mergers can take on relatively smooth, featureless distributions in kinematic spaces, but they can also exhibit complex structures with multiple overdensities in kinematic parameters (cf. Jean-Baptiste et al., 2017; Belokurov et al., 2023). Merger pairs, too, can also lead to well-separated stellar debris that is fairly easily decomposed into the individual dwarf galaxy components, but may also yield kinematic distributions in which the galaxies comprising the merger are indistinguishable, occupying the same regions of these kinematic spaces.

### v2

@@ III.3 Progenitors of GSE-like Debris
Generally, RA mergers are more recent ($t_{\mathrm{infall}}=5.9_{-2.0}^{+3.3}$ Gyr ago) and larger ($\log M_{\mathrm{dyn}}/\mathrm{M}_{\odot}=11.0^{+0.3}_{-0.1}$) than mergers contributing to RA pairs. Of those that contribute to RA pairs, the merger with the larger $f$ is typically more recent ($t_{\mathrm{infall}}=9.4^{+2.4}_{-2.7}$ versus $11.1^{+1.1}_{-2.4}$ Gyr ago) and has a slightly larger mass ($\log M_{\mathrm{dyn}}/\mathrm{M}_{\odot}=10.7^{+0.2}_{-0.4}$ versus $10.6^{+0.2}_{-0.7}$). This systematic shift in merger populations is particularly manifest in their SFHs, explored in Section V, but we first characterize the kinematic properties of these GSE analogs.
IV Kinematic Features
As summarized in Table 2, a number of GSE-like events are identified in TNG’s MW analogs. In general, RA debris is created by a single merger or—in a non-trivial fraction of cases—by pairs of mergers. A natural question that follows is whether or not these two scenarios can be distinguished using observable properties of the stellar halos. To this end, this section explores the kinematic properties of the stars in the RA mergers and pairs, such as their velocities, orbital energies and angular momenta. Section V will focus instead on stellar ages and abundances.
Figure 4 shows the distribution of spherical galactocentric velocities $v_{r}$ versus $v_{\phi}$ for the RA debris of four MW analogs. The star counts are indicated by the gray-scale histogram, while the contours enclose 68% of the stars from each merger: the single RA merger in red and the large-$f$ (small-$f$) merger of the RA pair in purple (pink). The four panels illustrate the kinematic diversity across different MWs in the TNG set. The top row of Figure 4 provides two examples of single RA mergers. In the left panel, the debris has a clear “sausage”-like velocity distribution, with a distinctive extension along $v_{r}$. In the right panel, the RA merger shows a greater population of stars on less eccentric orbits, with $v_{r}\sim 0$ and nonzero $v_{\phi}$, making the full extent of the debris potentially more difficult to distinguish from older contributions to the ex situ stellar halo that are expected to be nearly isotropic (Necib et al., 2019b).
Typically, single RA mergers have lower $v_{\phi}$ dispersion than RA pairs: the former have $\sigma_{\phi}=86^{+13}_{-18}~\mathrm{km\ s}^{-1}$ while the latter have $\sigma_{\phi}=95^{+14}_{-11}~\mathrm{km\ s}^{-1}$. However, because there is significant scatter in the azimuthal dispersion from halo to halo, the precise value of $\sigma_{\phi}$ does little to distinguish an individual RA merger from a pair. The other components of velocity are more comparable within their scatter. In the radial direction, single RA mergers exhibit a dispersion of $\sigma_{r}=159_{-27}^{+18}~\mathrm{km\ s}^{-1}$ versus the pairs’ $164_{-21}^{+29}~\mathrm{km\ s}^{-1}$. In the polar direction, RA mergers have $\sigma_{\theta}=90_{-20}^{+12}~\mathrm{km\ s}^{-1}$ versus the pairs’ $89_{-5}^{+12}~\mathrm{km\ s}^{-1}$. These trends (in addition to the trends in $\beta$ shown in the right panel of Figure 2) do indicate that debris from the RA pairs is slightly more isotropized, befitting of their earlier infall times (see Figure 3), though this is only evident in their $v_{\phi}$ values.
The debris from the same MWs is shown in Figure 5, where the distribution of orbital energy $E$ is plotted against $L_{z}$, the $z$-component of angular momentum. [^10] Because the shape of the distribution is more important than the numerical values of $E$ and $L_{z}$, the axes are rescaled. In particular, the energy is centered at $\bar{E}$, the median energy of the RA debris, and normalized by $\Delta E$, the difference between the 2nd and 98th percentiles. Similarly, though $L_{z}$ is still centered at zero, it is normalized by $\Delta L_{z}$, the difference between the 2nd and 98th percentile values of the RA debris’ $L_{z}$. As in Figure 4, the top row contrasts two examples of single RA mergers. The example on the right has more complex substructure, with debris deposited at distinct energies. On their own, distinct clumps in energy and angular momentum could potentially be misinterpreted as originating from multiple mergers—indeed, due to the inherent difficulties in effectively clustering stars in observational parameters, a standard approach is to accept this fact and intentionally over-fragment with clustering algorithms, so that multiple clusters of stars can later be associated to the same object (e.g. Koppelman et al., 2019; Yuan et al., 2020; Lövdal et al., 2022; Ou et al., 2023). These analyses must therefore be performed with great care, keeping in mind the potential complexities of stellar debris in these kinematic spaces.
Figure 6: SFHs for the ex situ stars of four MW analogs with RA debris. In each panel, the black histogram shows the total ex situ star formation, while the colored lines show the contribution from individual mergers, with single RA mergers in red, large-$f$ components of RA pairs in purple, and small-$f$ components of RA pairs in pink. The vertical dashed line shows the infall time, $t_{\mathrm{infall}}$, of each merger, and the dotted line shows the quenching time, $t_{90}$, by which 90% of stars have formed. The left column highlights cases in which star formation ceases shortly after the galaxies enter the virial radius of their host. In the case of the RA pair (which is given a marker consistent with previous figures), the periods of star formation of the two galaxies comprising the pair are well-separated in time, making them easier to distinguish from one another. The right column highlights examples where star formation continues long after the galaxies are accreted onto their hosts and where the RA pair does not show the separation seen in the left panel. Versions of this figure for each choice of RA debris are provided in C.
Figure 7: The quenching time $t_{90}$ relative to the infall time $t_{\mathrm{infall}}$, shown against the distance of first pericenter for each galaxy contributing to RA debris. All markers are consistent with Figure 3. Galaxies that continue forming stars after crossing the virial radius are below the gray $t_{90}=t_{\mathrm{infall}}{}$ line. While the quenching time is typically within 1 Gyr of the infall time, it is possible for star formation to continue well after the infalling galaxy crosses the virial radius. This is especially true of galaxies on orbits with large pericentric distances: in these cases, star formation can continue for many gigayears after infall.
Figure 8: Tinsley–Wallerstein diagram showing [Mg/Fe] abundance relative to [Fe/H] abundance for the same RA debris shown in Figure 6. Formatting is consistent with Figure 4 and Figure 5. A gold contour is also included for the in situ disk in each MW analog to guide the eye. RA debris is generally metal-poor and $\alpha$-enhanced relative to the disk. This chemistry can in some cases distinguish the two populations of stars comprising an RA pair, e.g. in the bottom left panel, which displays a clear $\alpha$-enhancement in the high-$f$ merger relative to the low-$f$ merger. However, this is not always the case; in the right column, the galaxies comprising the pair formed their stars at the same time, and the chemical abundances prove ineffective at separating the two. Versions of this figure for each choice of RA debris are provided in C.
Figure 9: Comparison of quantities relevant to star formation in each of the RA debris mergers. In each panel, the highlighted region runs from the 16th to the 84th percentile for the relevant quantity, with the scatter point at the 50th percentile (with markers consistent with previous figures). The Subfind ID of each MW analog is also provided. (Left:) Formation time, $t_{\mathrm{form}}$, of stars from each merger relative to the merger’s infall time, including a colored vertical bar indicating $t_{90}$, the time at which 90% of the galaxy’s stars had formed. Generally, star formation continues for some time prior to infall, at which point it ceases, and $t_{90}\sim t_{\mathrm{infall}}$. As RA pairs tend to have earlier infall times and smaller masses (both in $M_{\mathrm{dyn}}$ and $M_{\star}$), their star formation is brief and does not extend as far prior to $t_{\mathrm{infall}}$ as it does for RA mergers. Furthermore, RA mergers may hold on to their gas more effectively and continue forming stars many gigayears after $t_{\mathrm{infall}}$. (Center:) Abundance of iron relative to hydrogen. (Right:) Abundance of magnesium relative to iron. These right two panels show some systematic trends among the RA debris, with RA pairs more metal-poor and $\alpha$-enhanced than the single RA mergers, though these differences are less pronounced relative to the overall scatter in chemical abundances.
The bottom row of Figure 5 shows the RA pair examples in orbital energy and angular momentum space. In the left panel, the contributions from each merger are more easily distinguished from each other: only 12% of the high-$f$ merger stars fall within the pink contour, and 19% of the low-$f$ stars fall within the purple contour. Conversely, the right panel’s RA pair is not easily distinguished as two independent mergers, as 36% of the high-$f$ merger stars fall within the pink contour, and 50% of the low-$f$ fall within the purple contour. This overlap occurs even though the two mergers comprising the RA pair have infall times separated by more than 1 Gyr (Table 2) and have totally independent orbital initial conditions; their debris still settles into the same region of this kinematic space. There are hints that such overlap occurs in observational data, making it difficult to determine whether, e.g., structures like Sequoia, Arjuna, LMS-1, and I’itoi are truly independent structures or whether they should be considered part of the GSE debris (Koppelman et al., 2019; Massari et al., 2019; Malhan et al., 2021; Horta et al., 2023).
V Star Formation Histories
Chemodynamical space is particularly effective at separating various populations of ex situ material, as stars from different infalling dwarf galaxies will generically have different formation environments and therefore different chemical abundances. Chemical information can also be used to infer the ages of ex situ stars, which gives insight into the quenching time and overall SFH of the dwarf in which they originated. Because there are systematic shifts in the infall times of single mergers and pairs of mergers that contribute RA debris (see Section III.3), studying the SFH—and thence inferring the infall time of the dwarf galaxy progenitor(s)—is a promising route to distinguishing the two scenarios.
Infalling dwarfs are thought to stop forming stars quickly, as the MW’s circumgalactic medium strips them of their gas (Simpson et al., 2018; Akins et al., 2021; Engler et al., 2023). A close look at the satellites of MW-like hosts in TNG highlights the diversity of SFHs that can result from this process. In some cases—especially for the smallest, earliest-infalling satellites—star formation ceases soon after the satellite crosses the host’s virial radius. In other cases, the gas is not immediately stripped, and star formation continues for some time: for example, Engler et al. (2023) have shown that satellite quenching around MW analogs occurs 2.5 Gyr after infall for the median satellite in the TNG model and 7.5 Gyr after infall for the 84th percentile satellite. Figure 6 illustrates the diversity across four example MWs in our sample. The horizontal axis corresponds to the formation time of ex situ stars, $t_{\rm form}$. The black histogram is the total SFH, while the colored histograms are the contributions from the RA debris. The infall time of each merger, $t_{\mathrm{infall}}$, is indicated by a dashed vertical line of the same color. Each dotted vertical line corresponds to the quenching time, $t_{90}$, by which 90% of a dwarf’s stars have formed.
The left column of Figure 6 showcases galaxies in which star formation ceases quickly after infall, with $t_{\mathrm{infall}}{}\sim t_{90}$. Furthermore, the galaxies comprising the RA pair (bottom left panel) are well separated in terms of the formation times of their stars, meaning the SFHs of these two galaxies are sufficient to distinguish them from each other. The right column, however, highlights cases in which this is not the case. For the single merger, in the top right panel, star formation continues for some time following the dwarf’s infall, making it difficult to infer the $t_{\mathrm{infall}}$ from the SFH. For the RA pair, in the bottom right panel, the two galaxies form stars at very similar times, with overlapping SFHs that would make distinguishing the two stellar populations difficult on the basis of their age.
Whether or not star formation continues after $t_{\mathrm{infall}}$ is influenced by the galaxy’s orbit as it is accreted onto the MW analog. In particular, Figure 7 shows that continued star formation is correlated with large pericentric distances: the horizontal axis shows the distance of the infalling galaxy’s first pericenter, and the vertical axis shows the quenching time $t_{90}$ relative to the infall time $t_{\mathrm{infall}}$. Most galaxies—including those shown in Figure 6, save the top right panel—cease forming stars within 1 Gyr of their infall, but this is especially true for galaxies on orbits with small pericenters. In these cases, the galaxy falls directly toward the MW analog’s dense inner region, and star formation is prevented due to gas loss and tidal stripping. Galaxies with larger pericentric distances—such as ID 520885, shown in the top right panel of Figure 6, which has a pericentric distance of 77 kpc—are less affected by stripping and are able to continue forming stars for many gigayears after their infall (see Di Cintio et al., 2021).
Including chemical abundances, especially for $\alpha$-process elements, reveals more detailed information about the environment in which the stars formed. Figure 8 shows a Tinsley–Wallerstein diagram for the same halos as Figure 6, with formatting similar to previous figures: the grayscale histograms show the RA debris, with contours enclosing 68% of the stars from each merger. For reference, a gold contour encloses in situ disk stars, i.e. stars bound to the MW analog at the time of their formation that have orbital circularity $j_{z}/j_{\mathrm{circ}}>0.7$ (Scannapieco et al., 2009; Sotillo-Ramos et al., 2022). [^11] Here, $j_{z}$ is the $z$-component of the star’s specific angular momentum about the galactic center, and $j_{\mathrm{circ}}$ is the specific angular momentum of a circular orbit at the star’s current galactocentric distance.
While specific values of chemical abundances depend on the stellar physics model of TNG, [^12] The chemical abundances are in keeping with the current state-of-the-art, but they are based on metal yields that are rather uncertain (Pillepich et al., 2018). For example, it is known that TNG MW analogs are fairly $\alpha$-enriched relative to observations, see Naiman et al. (2018). Semenov et al. (2024, §2.3) provide approaches to correcting the metallicity distributions of TNG MWs, but in this work we instead focus on systematic shifts. relative shifts from one merger to another (and of mergers relative to the disk) still serve as an interesting comparison point. In the bottom row of Figure 8, the left panel shows the RA pair comprised of mergers with more distinct SFHs. Here, the two components have a clear separation in [Mg/Fe] relative to each other, with two sequences present in the histogram. In this pair, 10% of the high-$f$ merger stars fall within the pink contour and 2% of the low-$f$ merger stars fall within the purple contour. The bottom right panel shows the example RA pair with concurrent SFHs, and here the $\alpha$ abundance is not enough to clearly distinguish the two mergers from each other. In this pair, 54% of the high-$f$ merger stars fall within the pink contour and 39% of the low-$f$ merger stars fall within the purple contour.
Figure 9 summarizes the SFHs and chemical abundances across the sample: for each galaxy contributing to RA debris, a marker and horizontal line show the median and spread of the relevant quantity. The left panel shows the duration of star formation with respect to the infall time of the mergers, with an additional marker for the quenching time, $t_{90}$, at which 90% of the merger’s stars are formed. The duration of star formation is one suggestive distinction between RA mergers (in red) and RA pairs (in purple and pink), as the former have SFHs that extend well before their $t_{\mathrm{infall}}$, while the latter form their stars quickly, as they are generally accreted earlier in cosmic history (see Figure 3). The panel also reveals that $t_{90}$ generally serves as a good proxy for $t_{\mathrm{infall}}$; for most galaxies in this sample, the $t_{90}$ marker is close to the line through the origin that denotes the infall time. Notably, the exceptions to this trend are often single RA mergers, rather than the mergers of an RA pair; in terms of inferring $t_{\mathrm{infall}}$ from $t_{90}$, this further biases RA mergers toward later-inferred infall times.
The clear differences in SFH are less concretely reflected in chemical abundances in the TNG model, though two systematic shifts are still present. First, in [Fe/H], RA pairs are biased in the metal-poor direction relative to single RA mergers, indicative of the early epochs in which their stars formed. Second, in [Mg/Fe], the RA pairs are biased toward $\alpha$-enhancement, especially for the smaller-$f$ component of the pair (shown in pink). While these trends may not be sufficient to distinguish individual mergers from pairs (or galaxies comprising pairs from each other), the existence of the systematic biases between populations serves as a potential indicator for the merger histories of MW analogs. This is especially true if stellar ages can be robustly determined, as these trends are more apparent in SFH than in stellar chemistry.
VI Conclusion
This work presents a sample of 98 MW analogs in the TNG50 simulation of the IllustrisTNG suite. Each analog has its full accretion history identified with each ex situ star assigned to the merger that contributed it (Section II.3). This is the largest sample of MW analogs for which this decomposition has been performed, and it allows for unprecedented statistics in studying the assembly of the MW and the buildup of its stellar halo (Section III.1).
The MW analogs are chosen to match some properties of the Galaxy, namely its stellar mass and relative isolation (Section II.2), but certain aspects of the MW itself are rare across the sample. In particular, no analogs have an LMC-like satellite with mass $M_{\mathrm{dyn}}>10^{11}~\mathrm{M}_{\odot}$, indicating that the presence of the LMC distinguishes our Galaxy as an unlikely system. [^13] Recall that we make no requirement that the analogs have companion galaxies resembling the LMC, M31, or other Local Group objects. Additionally, while the MW analogs are more than 1 Mpc away from any cluster, their large-scale environments do not in general match that of our Galaxy. See Section II.2 for more details. Further, while 33% of MW analogs have stars with properties qualitatively similar to the GSE debris seen in our Galaxy—namely, highly radially biased velocities comprising a majority of the inner ex situ stars (Section III.2)—only 4% of MWs have debris as anisotropic as the GSE debris. This is in line with other studies that find the MW to be rare: see, e.g., Evans et al. (2020); Buch et al. (2024).
@@ VI Conclusion
Given the possibility of a two-merger GSE, the primary concern is in distinguishing it from a single-merger origin; however, it proves difficult for observable parameters to differentiate the two scenarios. Kinematics alone provide little conclusive evidence, as shown in Section IV. Single mergers can take on relatively smooth, featureless distributions in kinematic spaces, but they can also exhibit complex structures with multiple overdensities in kinematic parameters (see Jean-Baptiste et al., 2017; Belokurov et al., 2023). Merger pairs, too, can also lead to well-separated stellar debris that is fairly easily decomposed into the individual dwarf galaxy components, but may also yield kinematic distributions in which the galaxies comprising the merger are indistinguishable, occupying the same regions of these kinematic spaces.